trying to graph and label these systems of linear equations. State whether they are consistent, inconsistent dependent. If they are consistent, give the common solution.
x-y=4
5x-3y=8

Answer by rfer(16322)

x=y+4
5(y+4)-3y=8
5y+20-3y=8
2y=-12
y=-6
-----------------------
5x-3*-6=8
5x+18=8
5x=-10
x=-10/5
x=-2
--------------
-2+6=4
consistent
(-2,-6)